solve the elmination method -5x+6y=18
15x-18y=20

Answer by mananth(16946)

-5 x + 6 y = 18 .............1

15 x -18 y = 20 .............2
Eliminate y
multiply (1)by 3
Multiply (2) by 1
-15 x 18 y = 54
15 x -18 y = 20
Add the two equations
0 x = 74
Ther is no solution .
The lines are parallel.